### v6

@@ Appendix E Lemmas used in Laplace’s principle
Here, the first equality is because of Assumption 4.2 (B3) that $\nabla G\neq 0$ and the second equality comes from the co-area formula. $dH_{d-1}(v)$ is the (d1)(d-1) dimensional Hausdorff measure.
Now we fix $0<\tilde{u}<u$ and study the inner integral. We pick $\tilde{r}_{\tilde{u}}=\dfrac{(q+\mathcal{E}^{\tilde{u}}_{r}-\underline{\mathcal{E}_{\tilde{u}}})^{\mu}}{\eta}$. One can easily use Assumption 4.2 (C2) to verify the following facts:
  1. 1.
@@
For the proof of the first two facts, one can use the same method we used at the beginning of the proof of Lemma E.1 and details are omitted. For (3), one can prove it as follows:
@@
where the two inequalities are because of Assumption 4.2 (C2) and Lemma 4.6.
@@
where in the first and second inequalities, we used Assumption 4.2 (C1) that $\big{\|}v_{u}-v^{*}\big{\|}\leq\tau_{1}(u)$ and the fact that $\tau_{1}$ is an increasing function. Thus
@@
where in the equality, we used the co-area formula again and the definition of $\omega_{\alpha}$. Then combining with the fact that $\big{\|}\cdot\big{\|}_{\infty}\leq\big{\|}\cdot\big{\|}_{2}\leq\sqrt{d}\big{\|}\cdot\big{\|}_{2}$ finishes the proof.
Appendix F Proof of Laplace Principle: Lemma 4.7
Proof F.1.
By the definition of the consensus point $v_{\alpha}\left(\rho_{t}\right)$, one has
$\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$ $=\|\int v\cdot\dfrac{e^{-\alpha\mathcal{E}(v)}}{\|\omega_{\alpha}\|_{1}}\,d\rho_{t}(v)-v^{*}\|_{2}$
@@
$\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$ $\leq$ $\int\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$=$ $\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)+\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$+\int_{\{G\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v).$
For the first term, we can upper bound it using Lemma E.1:
$\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$ $\leq$ $\dfrac{\sqrt{d}(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v)$
@@
$+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v).$
For the second term, we can upper bound it using Lemma E.3:
$\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq$ $\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}$
$+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{t}(v)$
$+\sqrt{d}\tau_{1}(u).$
Finally, We leave the third term unchanged. Combining the estimates for the above three terms, we can finish the proof.
Appendix G The Complete Proof of Lemma 4.9
Proof G.1.
@@
$T_{1}(v)=-\lambda\left(v-v_{\alpha}\left(\rho_{t}\right)\right)\cdot\nabla\phi_{r}(v),$
$T_{2}(v)=\dfrac{\sigma^{2}}{2}\sum_{k=1}^{d}\left(v-v_{\alpha}\left(\rho_{t}\right)\right)_{k}^{2}\partial_{kk}\phi_{r}(v)$
@@
By the expression of $\nabla\phi_{r}$, one knows that $T_{3}\geq 0$ because of Assumption 4.2 (B1). Thus wone only has to find the lower bound of $T_{1}$ and $T_{2}$. The details of bounding them are exactly the same as [21] Proposition 2. Following the same steps, it turns out
@@
$=\int\left(T_{1}(v)+T_{2}(v)+T_{3}(v)\right)\,d\rho_{t}(v)\geq\int\left(T_{1}(v)+T_{2}(v)\right)\,d\rho_{t}(v)\geq-a\int\phi_{r}(v)\,d\rho_{t}(v).$
Then applying Gronwall’s inequality will finish the proof.
Appendix H Proof of Lemma 4.10
Proof H.1.
Let $B=\sup_{t\in[0,T]}\big{\|}v_{\alpha}\left(\rho_{t}\right)-v^{*}\big{\|}_{2}$ and $\tilde{B}=\sup_{t\in[0,T]}\mathcal{V}\left(\rho_{t}\right)$. Also, because of Assumption 4.2 (B2) that $G(v)\in C^{2}_{*}(\mathbb{R}^{d})$ and $G(v)\lesssim\big{\|}\nabla G(v)\big{\|}_{2}^{2}$, one can find some positive constant $\tilde{c}$ such that
@@
$\dfrac{d}{dt}\int G\,d\rho_{t}(v)$ $=$ $-\lambda\int\Bigl{\langle}v-v_{\alpha}\left(\rho_{t}\right),\nabla G\Bigr{\rangle}\,d\rho_{t}(v)+\dfrac{\sigma^{2}}{2}\int\sum_{k=1}^{d}\left(v-v_{\alpha}\left(\rho_{t}\right)\right)_{k}^{2}\partial_{kk}G\,d\rho_{t}(v)$
$-\dfrac{1}{\epsilon}\int\big{\|}\nabla G\big{\|}_{2}^{2}\,d\rho_{t}(v)$
$\leq$ $-\lambda\int\Bigl{\langle}v-v^{*},\nabla G\Bigr{\rangle}\,d\rho_{t}(v)-\lambda\int\Bigl{\langle}v^{*}-v_{\alpha}\left(\rho_{t}\right),\nabla G\Bigr{\rangle}\,d\rho_{t}(v)$
$+\sigma^{2}\int\sum_{k=1}^{d}\left((v-v^{*})_{k}^{2}+\left(v^{*}-v_{\alpha}\left(\rho_{t}\right)\right)_{k}^{2}\right)\partial_{kk}G\,d\rho_{t}(v)-\dfrac{1}{\epsilon}\int\big{\|}\nabla G\big{\|}_{2}^{2}\,d\rho_{t}(v).$
The first term is non-positive because of Assumption 4.2 (B1) and the second term can be bounded as follows
$-\lambda\int\Bigl{\langle}v^{*}-v_{\alpha}\left(\rho_{t}\right),\nabla G\Bigr{\rangle}\,d\rho_{t}(v)$ $\leq\lambda\int\big{\|}v^{*}-v_{\alpha}\left(\rho_{t}\right)\big{\|}\big{\|}\nabla G\big{\|}\,d\rho_{t}(v)$
@@
with initial condition $f(0)=\int G\,d\rho_{0}(v)$. By the comparison theorem, one knows that before $T$, $\int G\,d\rho_{t}(v)$ is dominated by $f$, i.e. $\int G\,d\rho_{t}(v)\leq f(t)$. And one has an explicit expression for $f$:
@@
$f(t)=\tilde{c}\epsilon D+\left(\int G\,d\rho_{0}(v)-\tilde{c}\epsilon D\right)e^{-(1/\tilde{c}\epsilon)t}\leq\tilde{c}\epsilon D+\left(\int G\,d\rho_{0}(v)-\tilde{c}\epsilon D\right)=\int G\,d\rho_{0}(v).$
@@
This completes the proof.
Appendix I Computational Details in the Proof of Theorem 4.1: Verification of the Assumptions in Lemma 4.7
Proof I.1.
@@
This verifies the assumptions in Lemma 4.7.
Appendix J Additional Numerical Example: Thomson’s Problem
We use Algorithm 2 with
[…]
and all the particle initially follow $V^{j}\sim\text{Unif}[-1,1]^{3k}$.
[…]
(70) $\text{relative error}=\frac{|\mathcal{E}(v_{\alpha}\left(\hat{\rho}\right))-\mathcal{E}(v^{*})|}{\mathcal{E}(v^{*})}$
[…]
Table 3: The result of Algorithm 2 on Thomson problem.
success rate relative error constraints value total steps
$k=2,(d=6)$ $100\%$ $4.4\times 10^{-3}$ $3.8\times 10^{-11}$ $382$
$k=3,(d=9)$ $100\%$ $9.9\times 10^{-3}$ $1.4\times 10^{-10}$ $407$
$k=8,(d=24)$ $100\%$ $1.78\times 10^{-2}$ $2.3\times 10^{-10}$ $567$
$k=15,(d=45)$ $100\%$ $1.57\times 10^{-2}$ $3.4\times 10^{-10}$ $895$
$k=56,(d=168)$ $97\%$ $1.44\times 10^{-2}$ $2.91\times 10^{-6}$ $1610$
$k=470,(d=1410)$ $93\%$ $1.95\times 10^{-2}$ $4.03\times 10^{-6}$ $1960$
(a) k=2
(b) k=3
(c) k=8
(d) k=15
(e) k=56
(f) k=470
Figure 7: Thomson Problem: the decay of the relative error over $100$ simulation and its mean.

### v8

@@ Appendix F Lemmas used in Laplace’s principle
Here, the first equality is because of Assumption 2 (B3) that $\nabla G\neq 0$ and the second equality comes from the co-area formula. $dH_{d-1}(v)$ is the (d1)(d-1) dimensional Hausdorff measure.
Now we fix $0<\tilde{u}<u$ and study the inner integral. We pick $\tilde{r}_{\tilde{u}}=\dfrac{(q+\mathcal{E}^{\tilde{u}}_{r}-\underline{\mathcal{E}_{\tilde{u}}})^{\mu}}{\eta}$. One can easily use Assumption 2 (C2) to verify the following facts:
  1. (1)
@@
For the proof of the first two facts, one can use the same method we used at the beginning of the proof of Lemma F.1 and details are omitted. For (3), one can prove it as follows:
@@
where the two inequalities are because of Assumption 2 (C2) and Lemma 4.4.
@@
where in the first and second inequalities, we used Assumption 2 (C1) that $\big\|v_{u}-v^{*}\big\|\leq\tau_{1}(u)$ and the fact that $\tau_{1}$ is an increasing function. Thus
@@
where in the equality, we used the co-area formula again and the definition of $\omega_{\alpha}$. Then combining with the fact that $\big\|\cdot\big\|_{\infty}\leq\big\|\cdot\big\|_{2}\leq\sqrt{d}\big\|\cdot\big\|_{2}$ finishes the proof. ∎
Appendix G Proof of Laplace Principle: Lemma 4.5
In this section, we present the proof of Lemma 4.5.
Proof.
By the definition of the consensus point $v_{\alpha}\big(\rho_{t}\big)$, one has
$\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$ $=\|\int v\cdot\dfrac{e^{-\alpha\mathcal{E}(v)}}{\|\omega_{\alpha}\|_{1}}\,d\rho_{t}(v)-v^{*}\|_{2}$
@@
$\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$
$\leq\int\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$=\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\quad+\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\quad+\int_{\{G\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v).$
For the first term, we can upper bound it using Lemma F.1:
$\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v)$
@@
$\quad+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v).$
For the second term, we can upper bound it using Lemma F.2:
$\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\quad+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{t}(v)$
$\quad+\sqrt{d}\tau_{1}(u).$
Finally, We leave the third term unchanged. Combining the estimates for the above three terms, we can finish the proof. ∎
Appendix H The Complete Proof of Lemma 4.6
Proof.
@@
$T_{1}(v)=-\lambda\left(v-v_{\alpha}\big(\rho_{t}\big)\right)\cdot\nabla\phi_{r}(v),$
$T_{2}(v)=\dfrac{\sigma^{2}}{2}\sum_{k=1}^{d}\left(v-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\partial_{kk}\phi_{r}(v)$
@@
By the expression of $\nabla\phi_{r}$, one knows that $T_{3}\geq 0$ because of Assumption 2 (B1). Thus wone only has to find the lower bound of $T_{1}$ and $T_{2}$. The details of bounding them are exactly the same as Proposition 2 [22]. Following the same steps, it turns out
@@
$=\int\left(T_{1}(v)+T_{2}(v)+T_{3}(v)\right)\,d\rho_{t}(v)$
$\geq\int\left(T_{1}(v)+T_{2}(v)\right)\,d\rho_{t}(v)\geq-a\int\phi_{r}(v)\,d\rho_{t}(v).$
Then applying Gronwall’s inequality will finish the proof. ∎
Appendix I Proof of Lemma 4.7
Proof.
Let $B=\sup_{t\in[0,T]}\big\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\big\|_{2}$ and $\tilde{B}=\sup_{t\in[0,T]}\mathcal{V}\big(\rho_{t}\big)$. Also, because of Assumption 2 (B2) that $G(v)\in C^{2}_{*}(\mathbb{R}^{d})$ and $G(v)\lesssim\big\|\nabla G(v)\big\|_{2}^{2}$, one can find some positive constant $\tilde{c}$ such that
@@
$\dfrac{d}{dt}\int G\,d\rho_{t}(v)$ $=$ $-\lambda\int\Bigl\langle v-v_{\alpha}\big(\rho_{t}\big),\nabla G\Bigr\rangle\,d\rho_{t}(v)$
$+\dfrac{\sigma^{2}}{2}\int\sum_{k=1}^{d}\left(v-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\partial_{kk}G\,d\rho_{t}(v)-\dfrac{1}{\epsilon}\int\big\|\nabla G\big\|_{2}^{2}\,d\rho_{t}(v)$
$\leq$ $-\lambda\int\Bigl\langle v-v^{*},\nabla G\Bigr\rangle\,d\rho_{t}(v)-\lambda\int\Bigl\langle v^{*}-v_{\alpha}\big(\rho_{t}\big),\nabla G\Bigr\rangle\,d\rho_{t}(v)$
$+\sigma^{2}\int\sum_{k=1}^{d}\left((v-v^{*})_{k}^{2}+\left(v^{*}-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\right)\partial_{kk}G\,d\rho_{t}(v)$
$-\dfrac{1}{\epsilon}\int\big\|\nabla G\big\|_{2}^{2}\,d\rho_{t}(v).$
The first term is non-positive because of Assumption 2 (B1) and the second term can be bounded as follows
$-\lambda\int\Bigl\langle v^{*}-v_{\alpha}\big(\rho_{t}\big),\nabla G\Bigr\rangle\,d\rho_{t}(v)$ $\leq\lambda\int\big\|v^{*}-v_{\alpha}\big(\rho_{t}\big)\big\|\big\|\nabla G\big\|\,d\rho_{t}(v)$
@@
with initial condition $f(0)=\int G\,d\rho_{0}(v)$. By the comparison theorem, one knows that before $T$, $\int G\,d\rho_{t}(v)$ is dominated by $f$, i.e., $\int G\,d\rho_{t}(v)\leq f(t)$. And one has an explicit expression for $f$:
@@
$f(t)$ $=\tilde{c}\epsilon D+\left(\int G\,d\rho_{0}(v)-\tilde{c}\epsilon D\right)e^{-(1/\tilde{c}\epsilon)t}$
$\leq\tilde{c}\epsilon D+\left(\int G\,d\rho_{0}(v)-\tilde{c}\epsilon D\right)=\int G\,d\rho_{0}(v).$
@@
This completes the proof. ∎
Appendix J Computational Details in the Proof of Theorem 4.1: Verification of the Assumptions in Lemma 4.5
Proof.
@@
This verifies the assumptions in Lemma 4.5. ∎